1. /2 points GrifEP7 6.E.12. At the low point in its swing, a pendulum bob with a mass of 0.7 kg has a velocity of 8 m/s (a) What is its kinetic energy at the low point? (b) Ignoring air resistance, how high will the bob swing above the low point before reversing direction?
Answer #1

a) kinetic energy at the low point

kinetic energy at the low point = 0.50 mv^2

kinetic energy at the low point = 0.50 * 0.7 * 8^2

kinetic energy at the low point = 22.4 J

b) let the height is h

m * g * h = kinetic energy

0.7 * 9.8 * h = 22.4

h = 3.26 m

the height is 3.26 m